OpenStudy (anonymous):

Please help me with this one question! A boat in calm seas travels in a straight line and ends the trip 22 km west and 53 km north of its original position. To the nearest tenth of a degree, find the direction of the trip. I am learning about vectors right now.

OpenStudy (tkhunny):

\(\theta\) degrees North of Due West. \(\tan(\theta) = 53/22\)